Mission Company is preparing its annual profit plan. As part of its analysis of the profitability of individual products, the controller estimates the amount of overhead that should be allocated to the individual product lines from the information provided below. (CMA based)

Wall Mirrors Specialty Windows
Units Produced 150 25
Material moves per product line 5 34
Direct labor hours per product line 750 850
Budgeted material handling costs: $336,000

Under a traditional costing system that allocates overhead on the basis of direct labor hours, the materials handling costs allocated to one unit of Wall Mirrors would be:

Multiple Choice

  • $1,000.

  • $1,050.

  • $4,200.

  • $16,000.
Answer #1

Answer:- $1050

Explanation:-

1 unit wall mirror = 336000/(750+850) = 210

= 210*(750)/150

= 1050